3x-6y= -15 and 18x+9y=20 are these lines parallel, perpendicular, or neither?
Answer by robertb(5830)

The first is the same as x-2y = -5, the second, the same as 2x + y = 20/9. Two lines of the form ax + by = c and bx - ay = d are perpendicular. Since these two lines are such, then they are perpendicular.